An aqueous solution of barium hydroxide is standardized by titration with a 0.199 M solution of hydroiodic acid.

If 19.1 mL of base are required to neutralize 18.2 mL of the acid, what is the molarity of the barium hydroxide solution?

______M barium hydroxide
Answer #1

Balanced chemical equation is:
2 HI + Ba(OH)2 ---> BaI2 + 2 H2O


Here:
M(HI)=0.199 M
V(HI)=18.2 mL
V(Ba(OH)2)=19.1 mL

According to balanced reaction:
1*number of mol of HI =2*number of mol of Ba(OH)2
1*M(HI)*V(HI) =2*M(Ba(OH)2)*V(Ba(OH)2)
1*0.199*18.2 = 2*M(Ba(OH)2)*19.1
M(Ba(OH)2) = 0.0948 M

Answer: 0.0948 M